OpenStudy (anonymous):

Help solve for: 2^x+3 – 2^x = 224

jimthompson5910 (jim_thompson5910):

\[\Large 2^{x+3} - 2^{x} = 224\] \[\Large 2^{x}*2^{3} - 2^{x} = 224\] \[\Large 2^{x}*8 - 2^{x} = 224\] \[\Large 8*2^{x} - 2^{x} = 224\] \[\Large 8z - z = 224 \ ... \ \text{Let } z = 2^x\] \[\Large 7z = 224\] I'll let you finish up.

jimthompson5910 (jim_thompson5910):

Keep in mind that your final answer will be of the form x = ... and not z = ...

OpenStudy (anonymous):

z= 224/7 z= 32 2^x = 32 2^5 = 32 x=5

OpenStudy (anonymous):

Thank you

jimthompson5910 (jim_thompson5910):

very good